d. 2 4. A researcher obtains M = 34 with SS = 190 for a sample of n = 10 girls and M = 29 with SS = 170 for a sample of n = 1 can you olease show the stems i got the answer C but the answer is B
Answer #1

Solution:

Given:

For girls:

Mg = 34

SSg = 190

ng = 10

For boys:

Mb = 29

SSb = 170

nb = 10

Thus t test statistic is given by:

M, - M \$(+ )

where

s – SS + 33 P nq + nb - 2

190 +170 10-10-2 “P

360

S_{p}^{2} =20

thus

M, - M \$(+ )

34 – 29 t = 20 x (1 + i)

t =- 20 x ()

t = 20 x ()

t = \frac{5}{\sqrt{ 4}}

الت |تا = 250

Thus correct option is B. الت |تا = 250